An elevator filled with passengers has a mass of 1,889 kilograms and moves upward at a speed of 9.4 meters/second for 1.8 seconds. Calculate the tension in the cable (in Newtons) supporting the elevator during this time.
Answer #1

acceleration

a = v/t = 9.4 / 1.8 = 5.22 m/s^2

mow reaction is given by

R = m ( g + a) = 1889* ("9.8 + 5.22)

R = 28377 N

=====

Comment before rate in case any doubt, will reply for sure.. goodluck